OpenStudy (anonymous):

simplify each: please explain how to do it rather than just giving me the answer

OpenStudy (anonymous):

ok what is it :)

OpenStudy (anonymous):

\[(1/y^2 - 1/x^2) / (x/y) +2+(y/x)\]

OpenStudy (primeralph):

@callie2240 Got this?

OpenStudy (primeralph):

Or whoever; I got to go.

OpenStudy (phi):

is it \[ \frac{ \frac{1}{y^2} - \frac{1}{x^2}}{\frac{x}{y}} + 2 + \frac{y}{x} \]

OpenStudy (anonymous):

@phi seems to have it :)

OpenStudy (phi):

or is it \[ \frac{ \frac{1}{y^2} - \frac{1}{x^2}}{\frac{x}{y} + 2 + \frac{y}{x} } \]

OpenStudy (anonymous):

the 2nd one

OpenStudy (phi):

if the 2nd way, you should put parens around the bottom so we know: (1/y2−1/x2)/ ( (x/y)+2+(y/x) )

OpenStudy (anonymous):

oh I meant to sorry

OpenStudy (phi):

The first step is get rid of the complicated bottom

OpenStudy (phi):

and to do that we should simplify it to a single fraction so for the time being, let's concentrate on \[ \frac{x}{y} + 2 + \frac{y }{x} \] any idea what the common denominator should be ?

OpenStudy (anonymous):

xy

OpenStudy (phi):

yes, Can you add up these terms, using xy as the common denominator ?

OpenStudy (anonymous):

yes hold on my computer is being stupid

OpenStudy (phi):

like this: \[ \frac{x}{y} \cdot \frac{x}{x} + 2 \cdot \frac{xy}{xy} + \frac {y}{x} \cdot \frac{y}{y}\]

OpenStudy (anonymous):

\[x^2/xy+2xy/xy+y^2/xy\]

OpenStudy (phi):

yes, which we can write as \[ \frac{x^2 + 2xy + y^2}{xy} \] we add the tops, and put that over the common denominator. now we use this idea: \[ \frac{a}{b} \cdot \frac{b}{a}= \frac{\cancel{a}}{\cancel{b}} \cdot \frac{\cancel{b}}{\cancel{a}}= 1\]

OpenStudy (phi):

in other words, if we multiply the complicated fraction's bottom by \[ \frac{xy}{x^2 + 2xy + y^2} \] it will turn into 1, and we can ignore it. of course, that means we have to multiply the top by the same thing

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

got it

OpenStudy (phi):

\[ \frac{ \frac{1}{y^2} - \frac{1}{x^2}}{\frac{x}{y} + 2 + \frac{y}{x} } = \frac{ \frac{1}{y^2} - \frac{1}{x^2}}{\frac{x^2+2xy+y^2}{xy}}\] now multiply top and bottom \[ \frac{ \frac{1}{y^2} - \frac{1}{x^2}\cdot \frac{xy}{x^2+2xy+y^2}}{\frac{x^2+2xy+y^2}{xy}\cdot \frac{xy}{x^2+2xy+y^2}} \]

OpenStudy (anonymous):

\[xy/(x^2y+2xy^2+y^4) - xy/(x^4+2x^2y+xy^2)\] all over 1

OpenStudy (phi):

there should be parens in that mess \[ \frac{ (\frac{1}{y^2} - \frac{1}{x^2})\cdot \frac{xy}{x^2+2xy+y^2}}{\frac{x^2+2xy+y^2}{xy}\cdot \frac{xy}{x^2+2xy+y^2}} \] the bottom is 1 \[ \frac{ (\frac{1}{y^2} - \frac{1}{x^2})\cdot \frac{xy}{x^2+2xy+y^2}}{1} \\= (\frac{1}{y^2} - \frac{1}{x^2})\cdot \frac{xy}{x^2+2xy+y^2}\]

OpenStudy (phi):

yes, but I would not multiply things. If there is any hope of simplifying it, we should factor it, and hope somethings will cancel. so instead, keep it like this \[ (\frac{1}{y^2} - \frac{1}{x^2})\cdot \frac{xy}{x^2+2xy+y^2} \] now subtract the two terms 1/y^2 - 1/x^2

OpenStudy (phi):

common denominator is x^2 y^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x^2/x^2y^2 - y^2/x^2y^2\]

OpenStudy (anonymous):

oh sorry forgot parenthesis

OpenStudy (anonymous):

hold on

OpenStudy (phi):

yes, or (x^2 - y^2)/ (x^2 y^2)

OpenStudy (anonymous):

can i multiply now?

OpenStudy (phi):

you do want to put the top over a common denominator \[ \frac{(x^2 -y^2)} {x^2 y^2} \]

OpenStudy (phi):

so far, we have \[ \frac{(x^2 -y^2)} {x^2 y^2} \cdot \frac{xy}{x^2+2xy+y^2} \] multiplying will give you the correct, (but unsimplified) answer. when you see a mess, think "factor" can you factor x^2 - y^2 (difference of squares has a factors nicely)

OpenStudy (anonymous):

(x-y)(x+y)

OpenStudy (phi):

can you factor the bottom x^2 + 2xy + y^2 hint: for a made up problem (unlike the real world) you might expect one of the factors up top to be in the bottom, so that things will cancel

OpenStudy (anonymous):

(x+y)(x+y)

OpenStudy (phi):

so we have \[ \frac{(x-y)(x+y)x y }{ x^2 y^2 (x+y)(x+y) }\] cancel time !

OpenStudy (anonymous):

wait what just happened? did we multiply?

OpenStudy (phi):

yes, top times top and bottom times bottom... not too tricky

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is this just the xy fraction on the left? or is it the whole thing?

OpenStudy (phi):

we got to \[ \frac{(x^2 -y^2)} {x^2 y^2} \cdot \frac{xy}{x^2+2xy+y^2} \] then you factored to get \[ \frac{(x -y)(x+y)} {x^2 y^2} \cdot \frac{xy}{(x +y)(x+y)} \] then (I should have let you do this step, because you like to multiply so much) : \[ \frac{ (x -y)(x+y) xy } { x^2y^2(x +y)(x+y)} \]

OpenStudy (anonymous):

ok and yeah it's fun

OpenStudy (phi):

now cancel terms that are in the top and bottom (that divide out to become 1)

OpenStudy (anonymous):

[(x-y)xy]/[x^2y^2(x+y)]

OpenStudy (phi):

did you forget how to simplify x/x^2 = x/(x*x) ?

OpenStudy (anonymous):

no I just forgot to do it

OpenStudy (anonymous):

x-y/ (xy)(x+y)

OpenStudy (phi):

yes though (x-y)/ (xy(x+y) ) is better x -y*stuff is different from (x-y)*stuff because the 2nd way means stuff multiplies both the x and the -y

OpenStudy (phi):

or divides in this case

OpenStudy (anonymous):

so is that it? am I done?

OpenStudy (phi):

yes

OpenStudy (anonymous):

yay! thank you! do you think you could help me with one more?

OpenStudy (phi):

if you go faster.

OpenStudy (anonymous):

i'll try..sorry